OpenStudy (anonymous):

Solve the system of equations. 4x + 7y = 6 and 6x + 5y = 20

OpenStudy (anonymous):

Someone answer this question pleeeeaaaassseee!!! :)

jhonyy9 (jhonyy9):

4x+7y=6 multiply by 3 6x+5y=20 multiply by -2 12x+21y=18 -12x-10y=-40 --------------- 0 11y=-22 so y= -22/11 = -2 y=-2 4x+7*(-2)=6 4x-14=6 4x=-20 x= -20/4 = -5 x=-5 y=-2

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

Thanks(:

jhonyy9 (jhonyy9):

yw good luck bye